# Accurate quantum-centric simulations of supramolecular interactions

### v1

@@ I Introduction
The accurate treatment of noncovalent interactions [muller2000noncovalent, puzzarini2020challenge] is extremely important in the biological, chemical, and pharmaceutical sciences [pichler2022important, aljoundi2020covalent]. Specifically, non-covalent interactions between hydrophobic species and hydrogen-bonded pairs play pivotal roles in a myriad of biological processes, ranging from protein folding [5, 6, 7, 8, 9]membrane assembly [10], cell signaling [11] and drug discovery  [Jhoti_Leach_2007, 12, 13, 14, 15]. The correct modeling of these interactions along with solvation plays a key role in understanding many chemical and biological processes  [hobza].
Traditionally, quantum mechanical [24, 25] methods have been used to study these systems to a high level of accuracy - so-called chemical accuracy (±1 kcal/mol from experiment). However, these calculations are quite expensive and approaches to accelerate these calculations continue to be explored using classical hardware. [Gao2024, Vogiatzis2017, Menczer2024, Ren2024, Jin2024, Schwenke2023, Delcey2023, Corzo2023, Datta2023, Zhai2021] Using the results from these calculations, force fields have been fine-tuned for wide use in molecular simulation studies of chemical and biological processes [RN49, NERENBERG2018129, RN51, RN52, RN53]. More recently, machine learning [RN50] methods built using accurate quantum chemical calculations on thousands of systems have appeared to study largely covalent interactions, but can be extended to non-covalent interactions at a good level of accuracy, but at a reduced computational cost. However, these latter methods build models that can can struggle to study diverse systems outside of the training set and can be subject to overfitting [27, 28].
Quantum computing based studies of these interactions, to a high level of accuracy and speed, would revolutionize our ability to understand complex processes like drug binding, but would also allow for the development of large synthetic datasets that could be used to build even better force fields and quantum machine learning models. However, to date, quantum hardware has struggled to address these problems. In this work we demonstrate that quantum-centric supercomputing (QCSC) [Alexeev2024] combined with the sample-based quantum diagonalization (SQD) approach [robledo2024chemistry] allows for the study of intermolecular interactions.
@@
The QCSC architecture enables scaling of computational capabilities, as exemplified by methods that use classical diagonalization in subspaces determined by quantum samples [robledo2024chemistry, kanno2023quantum]. These SQD methods use a quantum device to sample electronic configurations from a quantum circuit approximating the ground state of a molecular Hamiltonian, and use classical distributed HPC resources to post-process quantum measurements against known symmetries to obtain recovered configurations [robledo2024chemistry], as well as to solve the Schrödinger equation in the subspace spanned by the recovered configurations. The SQD method recently allowed us to address instances of the electronic structure problem with up to 36 spatial orbitals using up to 77 qubits [robledo2024chemistry]. These QCSC workflows produced significant improvements over simulations using quantum computers in isolation – which have in the last decade, used up to a handful of qubits with limited accuracies [Kandala2017, kandala2019error, google2020hartree, gao2021computational, gao2021applications, gocho2023excited, shirai2023computational, nam2020ground, khan2023chemically, obrien2023purification, zhao2023orbital, grimsley2019adaptive, christopoulou2024quantum, eddins2022doubling, motta2023quantum, castellanos2024quantum, liepuoniute2024simulation, weaving2024contextual, colless2018computation, dimitrov2023pushing, guo2024experimental, hempel2018quantum, huang2022variational, jones2024ground, kawashima2021optimizing, kirsopp2022quantum, leyton2023quantum, liang2024napa, liu2023performing, lolur2023reference, mccaskey2019quantum, rice2021quantum, santagati2018witnessing, shen2017quantum, smart2019quantum, yamamoto2022quantum, peruzzo2014variational, omalley2016scalable]. The QCSC paradigm coupled with SQD enables the study of problems heretofore out of reach of quantum computers including static correlation in iron-sulfur complexes [robledo2024chemistry] and well as dynamical correlation as exemplified in the intermolecular interactions studied herein.
@@
Past studies have reported the simulation of noncovalent interactions [loipersberger2023accurate, ollitrault2024estimation] using symmetry-adapted perturbation theory (SAPT). This method expresses the interaction energy through a perturbative treatment of the intermolecular potential [jeziorski1994perturbation, jeziorski1976first, patkowski2020recent], and requires the simulation of electronic structure of individual monomers on a quantum computer. In addition Anderson et al. demonstrated the possibility of simulations of coarse-grained intermolecular interactions on quantum computer as well. [anderson2022coarse] However, to date, predicting binding energies between monomers using the supramolecular approach, where the electronic structure of entire dimer need to be simulated on quantum hardware, has been an elusive target for quantum simulations, due to lack of accuracy and scale of conventional quantum approaches.
@@
Herein, we present the first quantum-centric simulation for the modeling of noncovalent hydrophilic and hydrophobic interactions with a supramolecular approach. We simulate the potential energy surfaces (PES) of the water dimer and the methane dimer. Our water dimer simulations use 27-qubit circuits, while the methane dimer simulations use 36- and 54-qubit circuits. To assess the accuracy of our quantum solutions, we compare them against heat-bath configuration interaction (HCI) [holmes2016heat, holmes2016efficient, smith2017cheap, sharma2017semistochastic] in the case of (16e,24o) calculations, complete active space configuration interaction for the (16e,12o) and (16e,16o) calculations, as well as coupled-cluster singles, doubles and perturbative triples (CCSD(T)) [bartlett2024perspective] performed for all of the studied instances. The latter is widely recognized as the gold standard for computing intermolecular interactions [carsky2010recent] to chemical accuracy. For the 27-qubit water dimer and the 36-qubit methane dimer simulations, we demonstrate that SQD energies agree with CASCI nearly exactly, while deviating from CCSD(T) within 1 kcal/mol in the equilibrium region of the PES. For the 54-qubit simulations of the methane dimer, we observe how the accuracy of the quantum solution can be systematically improved by increasing the number of sampled configurations.
@@ II.1 Classical benchmark
Metz et al  [metz2019molecular] and Li et al. [Li2009] demonstrated that CCSD(T)/aug-cc-pVQZ calculations closely reproduce the results of the CCSD(T)/complete basis set (CBS) limit for the methane dimer. Metz et al. [metz2019molecular] also demonstrated this for water dimer. All of our simulations are therefore done with the aug-cc-pVQZ basis set. We simulate the water and methane dimers with the active spaces listed in Table I.
We construct these active spaces using the atomic valence active space (AVAS) method [Sayfutyarova2017] as implemented in the PySCF 2.6.2 software package [sun2020recent, sun2018pyscf, sun2015libcint], and select active-space orbitals that overlap with the atomic orbitals (AOs) listed in column 3 of the table. The active-space orbitals of the water and methane dimers are shown in Fig. 1. Orbital visualization is performed with Pegamoid.[galvan_pegamoid]
@@
In each active space, we perform CCSD and CCSD(T) calculations with PySCF 2.6.2. For water and methane dimers we also perform CASCI(16e,12o) and CASCI(16e,16o) simulations, respectively, using PySCF 2.6.2. For the (16e,24o) active space of the methane dimer we perform HCI calculations with the SHCI-SCF 0.1 interface between PySCF 2.6.2 and DICE 1.0 [holmes2016heat, smith2017cheap, sharma2017semistochastic]. Further details of HCI calculations can be found in the Supplementary Information. Along with active-space simulations, we perform complete CCSD and CCSD(T) calculations with ORCA 5.0.4 [neese2022software]. The geometries of equilibrium structures of the water and methane dimer originate from works by Temelso et al [temelso2011benchmark] and by Rezac and Hobza [rezac2013describing], sourced through the BEGDB database [rezac2008quantum]. We describe the generation of PES geometries for water and methane dimers in the Supplementary Information.
@@ II.2 Quantum Computing
We prepare our wavefuntion guesses $|\Psi\rangle$, used to approximate the ground state of Eq. (3), from a truncated version of the local unitary cluster Jastrow (LUCJ) ansatz [motta2023bridging]
@@
where ${\hat{K}_{\mu}}=\Sigma_{pr,\sigma}K^{\mu}_{pr}\hat{a}_{p\sigma}^{\dagger}\hat{a}^{\phantom{\dagger}}_{r\sigma}$ are one-body operators, ${\hat{J}_{\mu}}=\Sigma_{pr,\sigma\tau}J^{\mu}_{p\sigma,r\tau}\hat{n}_{p\sigma}\hat{n}_{r\tau}$ are suitable (vide infra) density-density operators, and $|{\bf{x}}_{\mathrm{RHF}}\rangle$ is the restricted closed-shell Hartree-Fock (RHF) state. We use the Jordan-Wigner (JW) transformation [jordan1993paulische] to map the fermionic wavefunction Eq. (4) onto a qubit wavefunction that can be prepared executing a quantum circuit. The JW transformation maps the Fock space of fermions in $M$ spatial orbitals onto the Hilbert space of $2M$ qubits, where the basis state $|{\bf{x}}\rangle$ is parametrized by a bitstring ${\bf{x}}\in\{0,1\}^{2M}$ and represents an electronic configuration where the spin-orbital $p\sigma$ is occupied (empty) if $x_{p\sigma}=1$ ($x_{p\sigma}=0$). We prepare the wavefunction Eq. (4) by executing the following quantum circuit: a single layer of Pauli-X gates prepares the basis state $|{\bf{x}}_{\mathrm{RHF}}\rangle$, a Bogoliubov circuit [aleksandrowicz2019qiskit] (with linear depth, quadratic number of gates, and a 1D qubit connectivity) encodes each orbital rotation $e^{\pm\hat{K}_{\mu}}$, and a circuit of Pauli-ZZ rotations encodes each density-density interaction $e^{i\hat{J}_{\mu}}$. When $J^{\mu}$ is a dense matrix, Pauli-ZZ rotations are applied across all pair of qubits, requiring all-to-all qubit connectivity or a substantial overhead of swap gates. To mitigate these quantum hardware requirements LUCJ imposes a “locality” approximation, i.e., it assumes $J^{\mu}_{p\sigma,r\tau}=0$ for all pairs of spin-orbitals that are not mapped onto adjacent qubits under JW [motta2023bridging] (as a consequence, a circuit with constant depth and linear number of gates encodes each $e^{i\hat{J}_{\mu}}$ operator). Hence, the number of layers ($L-1$) in Eq. (4) is formally equal to 1.5. As the result the specific form of $|\Psi\rangle$ used in this work is expressed as $|\Psi\rangle=e^{-\hat{K}_{2}}e^{\hat{K}_{1}}e^{i\hat{J}_{1}}e^{-\hat{K}_{1}}|{\bf{x}}_{\mathrm{RHF}}\rangle$. We parametrize the LUCJ circuit based on amplitudes computed from classical restricted closed-shell CCSD within the given active space [robledo2024chemistry], yet a further quantum-classical parameter optimization could further improve the quality of the ground-state approximation. We produce the LUCJ circuits using the ffsim library [ffsim2024] interfaced with Qiskit 1.1.1 [aleksandrowicz2019qiskit, javadi2024quantum].
@@
The qubit layouts of the LUCJ circuits used for (16e,12o) water dimer, (16e,16o) methane dimer, and (16e,24o) methane dimer simulations are shown in Fig. 2a, 2b, and 2c, respectively. We execute these circuits on IBM’s 127-qubit Eagle devices ibm__\__cleveland and ibm__\__kyiv. In all our quantum computing experiments, we used gate (not measurement) twirling over random 2-qubit Clifford gates [wallman2016noise] and dynamical decoupling [viola1998dynamical, kofman2001universal, biercuk2009optimized, niu2022effects] – available through the SamplerV2 primitive of Qiskit’s runtime library – to mitigate quantum errors.
@@
in the form of bitstrings ${\bf{x}}\in\left\{0,1\right\}^{2M}$, each representing an electronic configuration (Slater determinants) distributed according to $\tilde{p}({\bf{x}})$. While on a noiseless device configurations are distributed according to $|\langle{\bf{x}}|\Psi\rangle|^{2}$, on a noisy device they follow a distribution $\tilde{p}({\bf{x}})\neq|\langle{\bf{x}}|\Psi\rangle|^{2}$. In particular, $\tilde{p}({\bf{x}})$ breaks particle-number conservation and returns configurations with incorrect particle number. We use a technique called self-consistent configuration recovery [robledo2024chemistry], executed on a classical computer, to restore particle-number conservation. The associated code is publicly available in the GitHub repository. [sqd_addon] Within each step of self-consistent recovery, we sample $K$ subsets (or batches) of $\tilde{\chi}$ labeled $\tilde{\chi}_{b}$ with $b=1\dots K$. Each batch defines – through a transformation [robledo2024chemistry] informed by an approximation to the ground-state occupation numbers $n_{p\sigma}$ – a subspace $S^{(b)}$ of dimension $d$, in which we project the many-electron Hamiltonian as [kanno2023quantum, nakagawa2023adapt, robledo2024chemistry]
@@
methane dimer (16e,24o) $200$ 5 $8.5$ $24.9\cdot 10^{7}$ 20, DICE 5
@@
We compute the ground-state eigenpairs of the Hamiltonians Eq. (6) using the iterative Davidson method on 10 CPUs with PySCF’s selected configuration interaction (SCI) solver for SQD (16e,12o) simulations of the water dimer and SQD (16e,16o) simulations of the methane dimer. We achieve parallelization across 10 CPUs with Ray 2.33.0 [moritz2017ray] where the eigenstate solver within each of the 10 batches is using 1 CPU. For SQD (16e,24o) simulations of the methane dimer, we utilize the SCI solver of DICE and 20 CPUs, where the eigenstate solver within each of the 5 batches is using 4 CPUs. Further parallelization is possible with the SCI solver of DICE, as was demonstrated previously [robledo2024chemistry]. The SQD (16e,12o) simulations of the water dimer and SQD (16e,16o) simulations of the methane dimer are done for the distances that are described in the Supplemental Information while SQD (16e,24o) simulations of the methane dimer are only done for 3.638 Å.

### v2

@@ I Introduction
The accurate treatment of noncovalent interactions [1, 2] is extremely important in the biological, chemical, and pharmaceutical sciences [3, 4]. Specifically, non-covalent interactions between hydrophobic species and hydrogen-bonded pairs play pivotal roles in a myriad of biological processes, ranging from protein folding [5, 6, 7, 8, 9]membrane assembly [10], cell signaling [11] and drug discovery  [12, 13, 14, 15, 16]. The correct modeling of these interactions along with solvation plays a key role in understanding many chemical and biological processes  [17].
Traditionally, quantum mechanical [18, 19] methods have been used to study these systems to a high level of accuracy - so-called chemical accuracy (±1 kcal/mol from experiment). However, these calculations are quite expensive and approaches to accelerate these calculations continue to be explored using classical hardware. [20, 21, 22, 23, 24, 25, 26, 27, 28, 29] Using the results from these calculations, force fields have been fine-tuned for wide use in molecular simulation studies of chemical and biological processes [30, 31, 32, 33, 34]. More recently, machine learning [35] methods built using accurate quantum chemical calculations on thousands of systems have appeared to study largely covalent interactions, but can be extended to non-covalent interactions at a good level of accuracy, but at a reduced computational cost. However, these latter methods build models that can can struggle to study diverse systems outside of the training set and can be subject to overfitting [36, 37].
Quantum computing based studies of these interactions, to a high level of accuracy and speed, would revolutionize our ability to understand complex processes like drug binding, but would also allow for the development of large synthetic datasets that could be used to build even better force fields and quantum machine learning models. However, to date, quantum hardware has struggled to address these problems. In this work we demonstrate that quantum-centric supercomputing (QCSC) [38] combined with the sample-based quantum diagonalization (SQD) approach [39] allows for the study of intermolecular interactions.
@@
The QCSC architecture enables scaling of computational capabilities, as exemplified by methods that use classical diagonalization in subspaces determined by quantum samples such as SQD [39] and QSCI [40]. The SQD method is developed based on QSCI. The SQD method use a quantum device to sample electronic configurations from a quantum circuit approximating the ground state of a molecular Hamiltonian, and use classical distributed HPC resources to post-process quantum measurements against known symmetries to obtain recovered configurations [39], as well as to solve the Schrödinger equation in the subspace spanned by the recovered configurations. The SQD method recently allowed us to address instances of the electronic structure problem with up to 36 spatial orbitals using up to 77 qubits [39]. The QCSC workflows produced significant improvements over simulations using quantum computers in isolation – which have in the last decade, used up to a handful of qubits with limited accuracies [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. The QCSC paradigm coupled with SQD enables the study of problems heretofore out of reach of quantum computers including static correlation in iron-sulfur complexes [39] and well as dynamical correlation as exemplified in the intermolecular interactions studied herein.
@@
Past studies have reported the simulation of noncovalent interactions [79, 80] using symmetry-adapted perturbation theory (SAPT). This method expresses the interaction energy through a perturbative treatment of the intermolecular potential [81, 82, 83], and requires the simulation of electronic structure of individual monomers on a quantum computer. In addition Anderson et al. demonstrated the possibility of simulations of coarse-grained intermolecular interactions on quantum computer as well. [84] However, to date, predicting binding energies between monomers using the supramolecular approach, where the electronic structure of entire dimer need to be simulated on quantum hardware, has been an elusive target for quantum simulations, due to lack of accuracy and scale of conventional quantum approaches.
@@
Herein, we present the first quantum-centric simulation for the modeling of noncovalent hydrophilic and hydrophobic interactions with a supramolecular approach. We simulate the potential energy surfaces (PES) of the water dimer and the methane dimer. Our water dimer simulations use 27-qubit circuits, while the methane dimer simulations use 36- and 54-qubit circuits. To assess the accuracy of our quantum solutions, we compare them against heat-bath configuration interaction (HCI) [85, 86, 87, 88] in the case of (16e,24o) calculations, complete active space configuration interaction for the (16e,12o) and (16e,16o) calculations, as well as coupled-cluster singles, doubles and perturbative triples (CCSD(T)) [89] performed for all of the studied instances. The latter is widely recognized as the gold standard for computing intermolecular interactions [90] to chemical accuracy. For the 27-qubit water dimer and the 36-qubit methane dimer simulations, we demonstrate that SQD energies agree with CASCI nearly exactly, while deviating from CCSD(T) within 1 kcal/mol in the equilibrium region of the PES. For the 54-qubit simulations of the methane dimer, we observe how the accuracy of the quantum solution can be systematically improved by increasing the number of sampled configurations.
@@ II.1 Classical benchmark
Metz et al  [91] and Li et al. [92] demonstrated that CCSD(T)/aug-cc-pVQZ calculations closely reproduce the results of the CCSD(T)/complete basis set (CBS) limit for the methane dimer. Metz et al. [91] also demonstrated this for water dimer. All of our simulations are therefore done with the aug-cc-pVQZ basis set. We simulate the water and methane dimers with the active spaces listed in Table I.
We construct these active spaces using the atomic valence active space (AVAS) method [93] as implemented in the PySCF 2.6.2 software package [94, 95, 96], and select active-space orbitals that overlap with the atomic orbitals (AOs) listed in column 3 of the table. The active-space orbitals of the water and methane dimers are shown in Fig. 1. Orbital visualization is performed with Pegamoid.[97]
@@
In each active space, we perform CCSD and CCSD(T) calculations with PySCF 2.6.2. For water and methane dimers we also perform CASCI(16e,12o) and CASCI(16e,16o) simulations, respectively, using PySCF 2.6.2. For the (16e,24o) active space of the methane dimer we perform HCI calculations with the SHCI-SCF 0.1 interface between PySCF 2.6.2 and DICE 1.0 [85, 87, 88]. Further details of HCI calculations can be found in the Supplementary Information. Along with active-space simulations, we perform complete CCSD and CCSD(T) calculations with ORCA 5.0.4 [98]. The geometries of equilibrium structures of the water and methane dimer originate from works by Temelso et al [99] and by Rezac and Hobza [100], sourced through the BEGDB database [101]. We describe the generation of PES geometries for water and methane dimers in the Supplementary Information.
@@ II.2 Quantum Computing
We prepare our wavefuntion guesses $|\Psi\rangle$, used to approximate the ground state of Eq. (3), from a truncated version of the local unitary cluster Jastrow (LUCJ) ansatz [102]
@@
where ${\hat{K}_{\mu}}=\Sigma_{pr,\sigma}K^{\mu}_{pr}\hat{a}_{p\sigma}^{\dagger}\hat{a}^{\phantom{\dagger}}_{r\sigma}$ are one-body operators, ${\hat{J}_{\mu}}=\Sigma_{pr,\sigma\tau}J^{\mu}_{p\sigma,r\tau}\hat{n}_{p\sigma}\hat{n}_{r\tau}$ are suitable (vide infra) density-density operators, and $|{\bf{x}}_{\mathrm{RHF}}\rangle$ is the restricted closed-shell Hartree-Fock (RHF) state. We use the Jordan-Wigner (JW) transformation [103] to map the fermionic wavefunction Eq. (4) onto a qubit wavefunction that can be prepared executing a quantum circuit. The JW transformation maps the Fock space of fermions in $M$ spatial orbitals onto the Hilbert space of $2M$ qubits, where the basis state $|{\bf{x}}\rangle$ is parametrized by a bitstring ${\bf{x}}\in\{0,1\}^{2M}$ and represents an electronic configuration where the spin-orbital $p\sigma$ is occupied (empty) if $x_{p\sigma}=1$ ($x_{p\sigma}=0$). We prepare the wavefunction Eq. (4) by executing the following quantum circuit: a single layer of Pauli-X gates prepares the basis state $|{\bf{x}}_{\mathrm{RHF}}\rangle$, a Bogoliubov circuit [104] (with linear depth, quadratic number of gates, and a 1D qubit connectivity) encodes each orbital rotation $e^{\pm\hat{K}_{\mu}}$, and a circuit of Pauli-ZZ rotations encodes each density-density interaction $e^{i\hat{J}_{\mu}}$. When $J^{\mu}$ is a dense matrix, Pauli-ZZ rotations are applied across all pair of qubits, requiring all-to-all qubit connectivity or a substantial overhead of swap gates. To mitigate these quantum hardware requirements LUCJ imposes a “locality” approximation, i.e., it assumes $J^{\mu}_{p\sigma,r\tau}=0$ for all pairs of spin-orbitals that are not mapped onto adjacent qubits under JW [102] (as a consequence, a circuit with constant depth and linear number of gates encodes each $e^{i\hat{J}_{\mu}}$ operator). Hence, the number of layers ($L-1$) in Eq. (4) is formally equal to 1.5. As the result the specific form of $|\Psi\rangle$ used in this work is expressed as $|\Psi\rangle=e^{-\hat{K}_{2}}e^{\hat{K}_{1}}e^{i\hat{J}_{1}}e^{-\hat{K}_{1}}|{\bf{x}}_{\mathrm{RHF}}\rangle$. We parametrize the LUCJ circuit based on amplitudes computed from classical restricted closed-shell CCSD within the given active space [39], yet a further quantum-classical parameter optimization could further improve the quality of the ground-state approximation. We produce the LUCJ circuits using the ffsim library [105] interfaced with Qiskit 1.1.1 [104, 106].
@@
The qubit layouts of the LUCJ circuits used for (16e,12o) water dimer, (16e,16o) methane dimer, and (16e,24o) methane dimer simulations are shown in Fig. 2a, 2b, and 2c, respectively. We execute these circuits on IBM’s 127-qubit Eagle devices ibm__\__cleveland and ibm__\__kyiv. In all our quantum computing experiments, we used gate (not measurement) twirling over random 2-qubit Clifford gates [107] and dynamical decoupling [108, 109, 110, 111] – available through the SamplerV2 primitive of Qiskit’s runtime library – to mitigate quantum errors.
@@
in the form of bitstrings ${\bf{x}}\in\left\{0,1\right\}^{2M}$, each representing an electronic configuration (Slater determinants) distributed according to $\tilde{p}({\bf{x}})$. While on a noiseless device configurations are distributed according to $|\langle{\bf{x}}|\Psi\rangle|^{2}$, on a noisy device they follow a distribution $\tilde{p}({\bf{x}})\neq|\langle{\bf{x}}|\Psi\rangle|^{2}$. In particular, $\tilde{p}({\bf{x}})$ breaks particle-number conservation and returns configurations with incorrect particle number. We use a technique called self-consistent configuration recovery [39], executed on a classical computer, to restore particle-number conservation. The associated code is publicly available in the GitHub repository. [112] Within each step of self-consistent recovery, we sample $K$ subsets (or batches) of $\tilde{\chi}$ labeled $\tilde{\chi}_{b}$ with $b=1\dots K$. Each batch defines – through a transformation [39] informed by an approximation to the ground-state occupation numbers $n_{p\sigma}$ – a subspace $S^{(b)}$ of dimension $d$, in which we project the many-electron Hamiltonian as [40, 113, 39]
@@
methane dimer (16e,24o) $300$ 4 $8.5$ $24.9\cdot 10^{7}$ 20, DICE 5
@@
We compute the ground-state eigenpairs of the Hamiltonians Eq. (6) using the iterative Davidson method on 10 CPUs with PySCF’s selected configuration interaction (SCI) solver for SQD (16e,12o) simulations of the water dimer and SQD (16e,16o) simulations of the methane dimer. We achieve parallelization across 10 CPUs with Ray 2.33.0 [114] where the eigenstate solver within each of the 10 batches is using 1 CPU. For SQD (16e,24o) simulations of the methane dimer, we utilize the SCI solver of DICE and 20 CPUs, where the eigenstate solver within each of the 5 batches is using 4 CPUs. Further parallelization is possible with the SCI solver of DICE, as was demonstrated previously [39]. The SQD (16e,12o) simulations of the water dimer and SQD (16e,16o) simulations of the methane dimer are done for the distances that are described in the Supplemental Information while SQD (16e,24o) simulations of the methane dimer are only done for 3.638 Å.